### v1

@@ F.2 Results
Figure 2 presents log-log plots of cumulative regret versus time for both algorithms. Each plot displays the mean regret computed over all 10 instances (with 10 runs each), along with shaded regions representing the standard deviation. The narrow confidence bands demonstrate the consistency of our results across different problem instances and random seeds. Note that the two plots use different x-axis representations: the EMC plot (Figure 2(a)) shows $\log(T)$ where each point represents the total regret after $T$ rounds of interaction, while the MvM plot (Figure 2(b)) shows $\log(t)$ with the cumulative regret trajectory at each round $t$.
@@
As shown in Figure 2(a), the algorithm exhibits consistent behavior, as evidenced by the nearly indistinguishable standard deviation bands. The fitted slope is approximately 0.694, approaching but not quite reaching the theoretical prediction of $2/3\approx 0.667$ from our $\tilde{O}(UKT^{2/3})$ regret bound. When tested for larger time horizons, the empirical slope approaches the theoretical value of $2/3$, confirming the asymptotic prediction of our analysis.
@@
As shown in Figure 2(b), the algorithm demonstrates excellent empirical performance with slopes ranging from 0.388 to 0.434 across different instances. All observed slopes fall strictly below the theoretical upper bound of 0.5, corresponding to our $\tilde{O}(U\sqrt{KT})$ regret bound. The variation in slopes across instances reflects the algorithm’s adaptive nature: instances with more challenging preference structures result in steeper regret growth, which corresponds to slower learning. The small standard deviation bands indicate that the algorithm’s performance is stable across multiple runs of the same instance. This gap between empirical performance and theoretical bounds suggests that our analysis may be conservative for typical problem instances.

### v2

@@ H.2 Results
Figure 2 presents log-log plots of cumulative regret versus time for both algorithms. Each plot displays the mean regret computed over all 10 instances (with 10 runs each), along with shaded regions representing the standard deviation. The narrow confidence bands demonstrate the consistency of our results across different problem instances and random seeds. Note that the two plots use different x-axis representations: the EMC plot (Figure˜2(a)) shows $\log(T)$ where each point represents the total regret after $T$ rounds of interaction, while the MvM plot (Figure˜2(b)) shows $\log(t)$ with the cumulative regret trajectory at each round $t$.
@@
As shown in Figure˜2(a), the algorithm exhibits consistent behavior, as evidenced by the nearly indistinguishable standard deviation bands. The fitted slope is approximately 0.694, approaching but not quite reaching the theoretical prediction of $2/3\approx 0.667$ from our $\tilde{O}(UKT^{2/3})$ regret bound. When tested for larger time horizons, the empirical slope approaches the theoretical value of $2/3$, confirming the asymptotic prediction of our analysis.
@@
As shown in Figure˜2(b), the algorithm demonstrates excellent empirical performance with slopes ranging from 0.388 to 0.434 across different instances. All observed slopes fall strictly below the theoretical upper bound of 0.5, corresponding to our $\tilde{O}(U\sqrt{KT})$ regret bound. The variation in slopes across instances reflects the algorithm’s adaptive nature: instances with more challenging preference structures result in steeper regret growth, which corresponds to slower learning. The small standard deviation bands indicate that the algorithm’s performance is stable across multiple runs of the same instance. This gap between empirical performance and theoretical bounds suggests that our analysis may be conservative for typical problem instances.
@@
Appendix I Bipartite Budgeted Matching Variant
In this appendix, we address a recommendation setting where each arm may serve at most one user.
I.1 Problem Formulation
We consider the same stochastic setting as in our main model $(T,U,K,(D_{u,k})_{u,k},(c_{k})_{k},B)$, but with a different notion of a valid matching.
In our main model, multiple users may be assigned to the same arm in a round, and the budget constraint applies to the set of distinct arms used. In the bipartite variant considered here, each user can be matched to at most one arm, and each arm can be matched to at most one user (one-to-one matching). For simplicity, we focus on the case of unit costs; the case of general costs remains an interesting problem for future work. In each round, the learner chooses a matrix $X\in\{0,1\}^{U\times K}$ satisfying
$\sum_{k}x_{u,k}\leq 1\ \forall u,\qquad\sum_{u}x_{u,k}\leq 1\ \forall k,\qquad\sum_{u,k}x_{u,k}\leq B,$
and receives an expected reward $V(\pi;\Theta)=\sum_{u,k}\Theta_{u,k}x_{u,k}$. Similarly to before, we denote by $\Pi^{\text{Bi}}$ the set of feasible matchings, where each matching correspond to a matrix that satisfy these constraints. The only change from our main model is the feasible set of matchings; the stochastic structure and regret definition remain unchanged.
I.2 Polynomial-Time Algorithm for the Offline Problem
We start our analysis by showing that the offline budgeted bipartite matching problem can be solved in polynomial time using a reduction to either the Hungarian algorithm or a min-cost flow formulation.
Proposition 18.
We can solve $\operatorname*{arg\,max}_{\pi\in\Pi^{\text{Bi}}}V(\pi;\Theta)$ in polynomial time.
Proof of Proposition˜18.
We divide our proof into two cases based on the relation between the budget and the number of users and arms.
Case 1: $B\geq\min(U,K)$.
In this case, the budget constraint is non-binding, and we can simply find the maximum weight matching in the bipartite graph formed by users and arms. Formally, we first pad the smaller side of the bipartite graph with dummy nodes to make it square. W.l.o.g., assume $U\leq K$; we add KUK-U dummy users with $\Theta_{u,k}=0$ for all arms $k$. We then apply the Hungarian algorithm to find the maximum weight matching in this square bipartite graph, which runs in $O(K^{3})$ time. Since the padded users do not contribute to the reward, the resulting matching is optimal for our original problem.
Case 2: $B<\min(U,K)$.
In this case, the budget constraint is binding, so we formulate the problem as a min-cost flow problem. We construct a flow network as follows:
  • Create a source node $s$ and a sink node $t$.
  • Create a node $A_{u}$ for each user $u$ and a node $B_{k}$ for each arm $k$.
  • Add edges from $s$ to each user node $A_{u}$ and from each arm node $B_{k}$ to $t$ with capacity 1 and cost 0.
  • For each user-arm pair $(u,k)$, add an edge from $A_{u}$ to $B_{k}$ with capacity 1 and cost $\Theta_{u,k}$
  • Node $s$ has a supply of $B$, and node $t$ has a demand of $B$.
  • Since the total supply at $s$ is $B$ and the total demand at $t$ is $B$, any valid flow of value $B$ corresponds to selecting exactly $B$ edges between users and arms.
By the Integrality Theorem of minimum cost flow, since all arc capacities and node supplies/demands are integers, there exists an integer-valued optimal flow $f^{*}$. In our network, since capacities are 1, this implies the flow on any edge $(A_{u},B_{k})$ is either 0 or 1. We construct the matching $X$ by setting $x_{u,k}=1$ if and only if the flow on edge $(A_{u},B_{k})$ is 1. The capacity constraints on edges $(s,A_{u})$ ensure $\sum_{k}x_{u,k}\leq 1$, and the capacity constraints on edges $(B_{k},t)$ ensure $\sum_{u}x_{u,k}\leq 1$. Finally, the total flow value of $B$ ensures $\sum_{u,k}x_{u,k}=B$.
The total cost of the flow is $\sum_{u,k}x_{u,k}(-\Theta_{u,k})=-\sum_{u,k}x_{u,k}\Theta_{u,k}$. Therefore, minimizing the cost is equivalent to maximizing the total reward $V(\pi;\Theta)$. Since min-cost flow can be solved in polynomial time, the offline bipartite budgeted matching problem is efficiently solvable. ∎
I.3 Online Algorithm
Since the offline optimization problem can be solved efficiently, we can employ standard Combinatorial MABs algorithms without relying on the single-peaked structure. Although we can adapt the algorithm from Corollary˜2 to this setting, we prefer to use the CombUCB1 algorithm from Kveton et al. [23] and its guarantees with our efficient offline solver as the oracle.
Corollary 19.
There exists an efficient online algorithm for the bipartite budgeted matching problem that achieves $\tilde{O}(\sqrt{T})$ regret.
Proof of Corollary˜19.
We can apply CombUCB1 using the offline solver from Proposition˜18 as the offline oracle. Since the reward function $V(\pi;\Theta)$ is linear in the matchings and the offline oracle is exact, the analysis of Kveton et al. [23] applies directly. The worst-case regret is bounded by $O(\sqrt{LK_{tot}T\ln T})$, where $L=UK$ is the number of total items (where item in this setting corresponds to choosing to match a user and item), and $K_{tot}=B$ is the maximum number of chosen items in a valid matching. Thus, the regret scales as $\tilde{O}(\sqrt{UKBT})$, and the per-round computational complexity is polynomial in $U,K$, and $B$. ∎
I.4 Lower Bound
We complement our algorithmic result with a matching lower bound, showing that the dependency on $U,K$, and $T$ is unavoidable.
Theorem 20.
For the bipartite budgeted matching problem with $K\geq U$, any online algorithm incurs an expected regret of $\Omega(\sqrt{UKT})$.
Proof of Theorem˜20.
Construct a hard instance where the set of arms $[K]$ is partitioned into $U$ disjoint subsets $S_{1},\dots,S_{U}$, each of size $|S_{u}|\geq 2$. User $u$ receives rewards only from arms in $S_{u}$. Specifically, for each user $u$, one arm in $S_{u}$ yields Bernoulli rewards with mean $1/2+\epsilon$, while others in $S_{u}$ have mean $1/2$. Arms outside $S_{u}$ yield 0. Since the subsets are disjoint and users are only compatible with their specific subsets, the global problem decomposes into $U$ independent multi-armed bandit instances. The constraints $\sum_{k}x_{u,k}\leq 1$ and $\sum_{u}x_{u,k}\leq 1$ are satisfied independently within each subset. Assuming sufficient global budget $B\geq U$, the total regret is the sum of regrets from $U$ independent bandits, each with $K/U$ arms over $T$ rounds. Using the standard minimax lower bound of $\Omega(\sqrt{K^{\prime}T})$ for $K^{\prime}$ arms:
$R(T)=\sum_{u=1}^{U}\Omega\left(\sqrt{\frac{K}{U}T}\right)=\Omega\left(U\sqrt{\frac{KT}{U}}\right)=\Omega\left(\sqrt{UKT}\right).$